# Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$

### v3

@@ 2. Description of the set of canonical forms
Proof. It is enough to show that $(x.\lambda)(e_{\gamma})$ can not be expressed as a linear combination of $\{(x.\lambda)(e_{\beta})\}_{\beta\succ\gamma}$. In other words, consider the orbit of the tangent action $\mathfrak{n}.\lambda$, and let $\mu\in\mathfrak{n}.\lambda$. It is enough to check that if all $\mu(e_{\beta})$ are known for $\beta\succ\gamma$, then the value of $\mu(e_{\gamma})$ can be arbitrary. We will present $\mu^{\prime}\in\mathfrak{n}.\lambda$ for which $\mu^{\prime}(e_{\beta})=\mu(e_{\beta})$ for all $\beta\succ\gamma$ and $\mu^{\prime}(e_{\gamma})=A$, where AA is a given value. Let $\mu^{\prime}=x.\lambda$, $x=\sum\limits_{\alpha\in\Phi^{+}}x_{\alpha}e_{\alpha}$. It follows from the proof of Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ that if $\delta\in\zeta_{D,\gamma}$ then $x_{\delta}$ is uniquely defined by the values $\{\mu(e_{\beta})\}_{\beta\succ\gamma}$. We define all other values $x_{\alpha}$, except $x_{\epsilon_{i}}$, as the corresponding values for $\mu$. We will find the values $x_{\epsilon_{i}}$ using induction on $i=\overline{1,k-1}$.
@@
Base. Let $i=1$, $A=\sum\limits_{\alpha\in\Phi^{+}}\lambda([e_{\delta_{1}},e_{\alpha}])x_{\alpha}$. Clearly, if $\lambda([e_{\delta_{1}},e_{\alpha}])\neq 0$ then $\alpha=\epsilon_{1}$ or $\alpha\in\zeta_{D,\gamma}$. Hence, all values $x_{\alpha}$ with nonzero $\lambda([e_{\delta_{1}},e_{\alpha}])$ in the right hand side, except $x_{\epsilon_{1}}$, are already defined. Since that $\lambda([e_{\delta_{1}},e_{\epsilon_{1}}])\neq 0$, the value of $x_{\epsilon_{1}}$ is uniquely determined, too. The only $\mu^{\prime}(e_{\beta}),\beta\succeq\gamma$ that can still depend on $x_{\epsilon_{1}}$ is $\mu^{\prime}(e_{\delta_{2}})$. If $k=1$, then nothing else depends on $x_{\epsilon_{1}}$.
@@
Step. Let the value of $x_{\epsilon_{i-1}}$ be found. Consider $\mu(e_{\delta_{i}})=\sum\limits_{\alpha\in\Phi^{+}}\lambda([e_{\delta_{i}},e_{\alpha}])x_{\alpha}$. On the right side, there can only be $x_{\epsilon_{i-1}},x_{\epsilon_{i}}$ and $x_{\alpha}$, for which $\alpha\in\zeta_{D,\gamma}$. The value of $x_{\epsilon_{i-1}}$ was determined in the previous step. That is, on the right side, only $x_{\epsilon_{i}}$ is undefined, and it is definitely there. From this condition, the value of $x_{\epsilon_{1}}$ is uniquely determined. The only $\mu^{\prime}(e_{\beta}),\beta\succ\gamma$ that still depends on $x_{\epsilon_{i}}$ is $\mu^{\prime}(e_{\delta_{i+1}})$.
@@
Case $i=k$. Consider $\mu(e_{\delta_{k}})=\sum\limits_{\alpha\in\Phi^{+}}\lambda([e_{\delta_{k}},e_{\alpha}])x_{\alpha}$. The right side definitely has $x_{\epsilon_{k}}$ and possibly $x_{\epsilon_{k-1}}$. The value of $x_{\varepsilon_{k-1}}$ was determined in the previous step (or it is not there at all if $k=1$). All other values $x_{\alpha}$ are also defined. Therefore, the value of $x_{\epsilon_{k}}$ is uniquely determined. Moreover, no other $\mu^{\prime}(e_{\beta}),\beta\succeq\gamma$ depend on $x_{\epsilon_{k}}$, and its value no longer affects anything else.
Thus, all values $x_{\alpha}$ were calculated and the required $\mu^{\prime}$ was presented. $\square$
@@
Also note that the conditions of Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ do not depend on the coordinates of $\lambda$, but only on its support. Therefore, it is enough to check these conditions only for one form with a given support, for example, we may assume that all nonzero coordinates equal 11.
@@
The conditions of Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ and Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ are formulated so that they can be checked using programs. Below we present algorithms for checking the necessary and sufficient conditions for a support belonging to the set $S$.
@@
An incomplete enumeration of supports is performed. For each support, the necessary (respectively, sufficient) condition is checked, and then the next support to be checked is built based on it. The enumeration is based on the following fact. Let $D$ be a support, $\alpha$ a positive root that is not in $D$ and is smaller than all roots from $D$. Then if $D$ does not satisfy the necessary (respectively, sufficient) condition, then $D\cup\{\alpha\}$ also does not satisfy the necessary (respectively, sufficient) condition. Indeed, the necessary (respectively, sufficient) condition for the support does not hold when there exists a root $\gamma$ in the support that satisfies certain conditions. These conditions depend only on $\beta\in D$ such that $\beta\succeq\gamma$. Neither the set $\zeta_{D,\gamma}$, nor the values $\lambda([e_{\alpha},e_{\beta}])$, where $\alpha$ or $\beta\succeq\gamma$, depend on the roots in the support that are smaller than $\gamma$.
@@
The enumeration is performed according to the following rules. First, the support $\{2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4}\}$ is checked for compliance with the necessary (respectively, sufficient) condition. Further, let us assume that we have checked the support $D$ for compliance with the necessary (respectively, sufficient) condition. Let $\delta$ be the minimal root in $D$. First, let us consider the case when $\delta\neq\alpha_{4}$. If $D$ passed the check, we add to $D$ the positive root closest to $\delta$ from below. If $D$ did not pass the check, we remove $\delta$ from $D$ and add the positive root closest to $\delta$ from below. Then we check the new support for compliance with the necessary (respectively, sufficient) condition. Second, if $\delta=\alpha_{4}$, then we remove $\delta$, remove the new minimal root $\beta$ in the support (if exists) and add the root closest to $\beta$ from below. If $D=\{\alpha_{4}\}$, then the enumeration ends here. In the end, we include the empty set in the list of supports, because it definitely satisfies the necessary (respectively, sufficient) condition. Thus, we will enumerate all supports that potentially satisfy the necessary (respectively, sufficient) condition, and we will not check supports that definitely do not satisfy these conditions.
@@
Let us clarify that if $D$ is the support of a form $\lambda$, then the condition $\lambda([e_{\alpha},e_{\beta}])\neq 0$ is equivalent to $\alpha+\beta\in D$. Checking this condition can be easily implemented. As it was already mentioned, the necessary (respectively, sufficient) condition needs to be checked for all roots of the support $D$. It is obvious that it is enough to check it only for the root $\gamma$ that we just added to the support, because for the rest of the roots we checked the condition before. The first thing to do is to build the set $\zeta_{D,\gamma}$. It can be built as follows. We launch a cycle and go through the rows that are numbered by roots $\alpha\succ\gamma$. If there is only one non-zero element $\lambda([e_{\alpha},e_{\delta}])$ in the row $\alpha$ such that $\delta\notin\zeta_{D,\gamma}$, then we add $\delta$ to $\zeta_{D,\gamma}$. If, after performing the cycle, the number of roots in $\zeta_{D,\gamma}$ has increased, we go through it again. The process stops when the cardinality of $\zeta_{D,\gamma}$ stabilizes. Thus, we have built $\zeta_{D,\gamma}$.
Checking the sufficient condition for the support $D$ is trivial. We go through the row numbered $\gamma$ and check the condition of Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$. The necessary condition is checked as follows. Let us describe the $i$-th step of the check. We go through the row numbered $\delta_{i}$ (recall that $\delta_{1}=\gamma$). If we meet a non-zero element $\lambda([e_{\delta_{i}},e_{\varepsilon}])$ such that $\varepsilon\notin\zeta_{D,\gamma}$ and $\varepsilon\neq\varepsilon_{i-1}$, we check the column numbered $\varepsilon$, starting from $\gamma$ and higher. If there are no non-zero elements in this part of the column, except for $\lambda([e_{\delta_{i}},e_{\varepsilon}])$, then the necessary condition is not satisfied. If this is not the case, but only one non-zero element was found in the row (possibly except for $\lambda([e_{\delta_{i}},e_{\varepsilon_{i-1}}])$) of the form $\lambda([e_{\delta_{i}},e_{\varepsilon}])$ such that $\varepsilon\notin\zeta_{D,\gamma}$, and at the same time, only one non-zero element was found in the corresponding part of the column, except for ours, of the form $\lambda([e_{\delta},e_{\varepsilon}])$, then $\varepsilon=\varepsilon_{i}$, and $\delta=\delta_{i+1}$. And then we check the condition already for the column numbered $\delta_{i+1}$. In fact, our program allows checking the necessary condition only up to the second step. And if the necessary condition is still not violated by that moment, it considers that the check has passed. But this hardly complicates the final calculation.
@@
The results of the check are as follows: 878 supports satisfy the sufficient condition, while 911 satisfy the necessary condition, so 33 supports remain to be checked manually. Here they are:
@@
Linear forms with support 5 actually don’t satisfy the necessary condition, but checking this requires three steps, while the algorithm checks only two steps, see above. It is clear that linear forms with the supports numbered 2, 6, 12, 14, 16, 18, 19, 20 and 21 do not satisfy condition (1), i.e., such forms do not belong to $S$. Indeed, it follows immediately from Table 1 as in Example Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$. In other cases, it is necessary to explicitly write out the coordinates of the tangent action. It turns out that linear forms with the supports numbered 25 and 33 lie in $S$ without restrictions on the coordinates. Linear forms with supports numbered 30, 31 and 32 lie in $S$ if and only if $\lambda_{\alpha_{1}+\alpha_{2}}\lambda^{2}_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}=\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}\lambda^{2}_{\alpha_{2}+\alpha_{3}+\alpha_{4}}$. We obtain 883 supports whose forms belong to $S$, 3 of which require a restriction on the coordinates of forms. Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ is proved.
Example 2.5. Consider a form $\lambda$ with support 33. We will show that $\lambda\in S$. Let $\mu\in\mathfrak{n}.\lambda,~\mu=x.\lambda,~x=\sum\nolimits_{\alpha\in\Phi^{+}}x_{\alpha}e_{\alpha},~\lambda=\sum\nolimits_{\alpha\in\Phi^{+}}\lambda_{\alpha}e_{\alpha},~\mu=\sum\nolimits_{\alpha\in\Phi^{+}}\mu_{\alpha}e_{\alpha}$. Let us explicitly write down the coordinates we are interested in. We will denote the structural constants in the expression $Ce_{\alpha+\beta}=[e_{\alpha},e_{\beta}]$ by $C_{\alpha,\beta}$. Their values can be found in [KP23]. Then
@@
Let us also consider the support numbered 25. All roots, except $\alpha_{2}$, obviously satisfy condition (1). The coordinate $\mu_{\alpha_{2}}$ is equal to $x_{\alpha_{3}}\lambda_{\alpha_{2}+\alpha_{3}}$, so if the polynomial $x_{\alpha_{3}}$ can be expressed as a linear combination of $\{\mu_{\beta}\}_{\beta\succ\alpha_{2}}$ then $\alpha_{2}$ satisfies condition (1). That is exactly the case:
@@
Finally, let us consider forms with the supports 30, 31 and 32. Let the support contain roots $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4}$. Let us check whether condition (1) holds for the root $\alpha_{2}$. Let us note that
@@
We see that if $\beta\succ\alpha_{2}$ and $x_{\alpha_{3}+\alpha_{4}}$ enters $\mu_{\beta}$ then $\beta=\alpha_{2}+\alpha_{3}$; similarly, if $\beta\succ\alpha_{2}$ and $x_{\alpha_{1}}$ enters $\mu_{\beta}$ then $\beta=\alpha_{2}+2\alpha_{3}$. Hence, the coordinate $\mu_{\alpha_{2}}$ should be a linear combination of $\mu_{\alpha_{2}+2\alpha_{3}}$ and $\mu_{\alpha_{2}+\alpha_{3}}$. It is easy to check that this is true if and only if $\lambda_{\alpha_{1}+\alpha_{2}}\lambda^{2}_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}=\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}\lambda^{2}_{\alpha_{2}+\alpha_{3}+\alpha_{4}}$.
Let us check whether condition (1) holds for the root $\alpha_{4}$. Note that
@@
Arguing as above we obtain that the coordinate $\mu_{\alpha_{4}}$ should be a linear combination of $\mu_{\alpha_{1}}$ and $\mu_{\alpha_{3}+\alpha_{4}}$. It turns out that this is true if and only if $\lambda_{\alpha_{1}+\alpha_{2}}\lambda^{2}_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}=\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}\lambda^{2}_{\alpha_{2}+\alpha_{3}+\alpha_{4}}$.

### v4

@@ 2. Description of the set of canonical forms
Proof. It is enough to show that $(x.\lambda)(e_{\gamma})$ cannot be expressed as a linear combination of $\{(x.\lambda)(e_{\beta})\}_{\beta\succ\gamma}$. In other words, consider the orbit of the tangent action $\mathfrak{n}.\lambda$, and let $\mu\in\mathfrak{n}.\lambda$. It is enough to check that if all values $\mu(e_{\beta})$ are known for $\beta\succ\gamma$, then the value of $\mu(e_{\gamma})$ can be arbitrary. We will construct $\mu^{\prime}\in\mathfrak{n}.\lambda$ such that $\mu^{\prime}(e_{\beta})=\mu(e_{\beta})$ for all $\beta\succ\gamma$ and $\mu^{\prime}(e_{\gamma})=A$, where AA is an arbitrary given value. Let $\mu^{\prime}=x.\lambda$, $x=\sum\limits_{\alpha\in\Phi^{+}}x_{\alpha}e_{\alpha}$. It follows from the proof of Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ that if $\delta\in\zeta_{D,\gamma}$, then $x_{\delta}$ is uniquely defined by the values $\{\mu(e_{\beta})\}_{\beta\succ\gamma}$. We define all other values $x_{\alpha}$, except $x_{\epsilon_{i}}$, to be the same as for $\mu$. We will determine the values $x_{\epsilon_{i}}$ by induction on $i=\overline{1,k-1}$.
@@
Base case. Let $i=1$, $A=\sum\limits_{\alpha\in\Phi^{+}}\lambda([e_{\delta_{1}},e_{\alpha}])x_{\alpha}$. Clearly, if $\lambda([e_{\delta_{1}},e_{\alpha}])\neq 0$, then either $\alpha=\epsilon_{1}$ or $\alpha\in\zeta_{D,\gamma}$. Hence, in the right-hand side, all values $x_{\alpha}$ with nonzero $\lambda([e_{\delta_{1}},e_{\alpha}])$ except possibly $x_{\epsilon_{1}}$, are already defined. Since $\lambda([e_{\delta_{1}},e_{\epsilon_{1}}])\neq 0$, the value of $x_{\epsilon_{1}}$ is uniquely determined as well. After this, the only expression $\mu^{\prime}(e_{\beta}),\beta\succeq\gamma$ that may still depend on $x_{\epsilon_{1}}$ is $\mu^{\prime}(e_{\delta_{2}})$. If $k=1$, then nothing else depends on $x_{\epsilon_{1}}$.
@@
Inductive step. Suppose the value $x_{\epsilon_{i-1}}$ has been found. Consider $\mu(e_{\delta_{i}})=\sum\limits_{\alpha\in\Phi^{+}}\lambda([e_{\delta_{i}},e_{\alpha}])x_{\alpha}$. On the right-hand side, only $x_{\epsilon_{i-1}},x_{\epsilon_{i}}$ and $x_{\alpha}$, for which $\alpha\in\zeta_{D,\gamma}$, can appear with non-zero coefficients. The value $x_{\epsilon_{i-1}}$ was determined in the previous step. Thus, on the right-hand side, only $x_{\epsilon_{i}}$ may be undefined, and it definitely appears. From this equation, the value $x_{\epsilon_{1}}$ is uniquely determined. The only $\mu^{\prime}(e_{\beta}),\beta\succ\gamma$ that may still depend on $x_{\epsilon_{i}}$ is $\mu^{\prime}(e_{\delta_{i+1}})$.
@@
Case $i=k$. Consider $\mu(e_{\delta_{k}})=\sum\limits_{\alpha\in\Phi^{+}}\lambda([e_{\delta_{k}},e_{\alpha}])x_{\alpha}$. The right-hand side definitely contains $x_{\epsilon_{k}}$ and possibly $x_{\epsilon_{k-1}}$. The value $x_{\varepsilon_{k-1}}$ was determined in the previous step (or does not appear if $k=1$). All other $x_{\alpha}$ are also defined. Therefore, the value $x_{\epsilon_{k}}$ is uniquely determined. Moreover, no other $\mu^{\prime}(e_{\beta}),\beta\succeq\gamma$ depends on $x_{\epsilon_{k}}$, and its value no longer affects anything else.
Thus, we have computed all coefficients $x_{\alpha}$ and constructed the required $\mu^{\prime}$. $\square$
@@
Also note that the conditions of Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ do not depend on the specific coordinates of $\lambda$, but only on its support. Therefore, it is sufficient to check these conditions for just one form with a given support; for example, we may assume that all its nonzero coordinates are equal to 11.
@@
The conditions of Propositions Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ and Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ are formulated so that they can be checked algorithmically. Below we present algorithms for checking the necessary and sufficient conditions for a support to belong to the set SS."
@@
An incomplete enumeration of supports is performed. For each support, the necessary (respectively, sufficient) condition is checked, and the next support to be checked is then generated based on the result. The enumeration relies on the following observation. Let $D$ be a support, and let $\alpha$ be a positive root not in $D$ that is smaller than all roots in $D$. If $D$ does not satisfy the necessary (respectively, sufficient) condition, then $D\cup{\alpha}$ also fails to satisfy the necessary (respectively, sufficient) condition. Indeed, the necessary (respectively, sufficient) condition for a support fails when there exists a root $\gamma$ in the support satisfying certain conditions. These conditions depend only on those $\beta\in D$ such that $\beta\succeq\gamma$. Neither the set $\zeta_{D,\gamma}$ nor the values $\lambda([e_{\alpha},e_{\beta}])$ (where $\alpha$ or $\beta\succeq\gamma$) depend on the roots in the support that are smaller than $\gamma$.
@@
The enumeration proceeds according to the following rules. First, the support $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4}$} is checked against the necessary (respectively, sufficient) condition. Next, suppose we have checked a support $D$ against the necessary (respectively, sufficient) condition. Let $\delta$ be the minimal root in $D$. First, let us consider the case if $\delta\neq\alpha_{4}$. If $D$ passes the check, we add to $D$ the positive root immediately preceding $\delta$ (with respect to the order $\succ$). If $D$ fails the check, we remove $\delta$ from $D$ and add the positive root immediately preceding $\delta$. Then we check the new support against the necessary (respectively, sufficient) condition. Second, if $\delta=\alpha_{4}$, then we remove $\delta$, remove the new minimal root $\beta$ in the support (if it exists) and add the root closest to $\beta$ from below. If $D=\{\alpha_{4}\}$, then the enumeration ends here. Finally, we include the empty set in the list of supports, since it trivially satisfies the necessary (respectively, sufficient) condition. In this way, we enumerate all supports that could potentially satisfy the necessary (respectively, sufficient) condition, while avoiding checks for supports that certainly do not satisfy these conditions.
@@
Note that if $D$ is the support of a form $\lambda$, then the condition $\lambda([e_{\alpha},e_{\beta}])\neq 0$ is equivalent to $\alpha+\beta\in D$. Checking this condition is easy to implement. As mentioned earlier, the necessary (respectively, sufficient) condition must be checked for all roots in the support $D$. It is clear that it suffices to check it only for the root $\gamma$ that was most recently added to the support, because for all other roots the condition has already been verified. The first step is to construct the set $\zeta_{D,\gamma}$. This can be done as follows. We run a loop over rows indexed by roots $\alpha\succ\gamma$. If in a row $\alpha$ there is exactly one non-zero element $\lambda([e_{\alpha},e_{\delta}])$ such that $\delta\notin\zeta_{D,\gamma}$, then we add $\delta$ to $\zeta_{D,\gamma}$. After completing a full pass through the rows, if the size of $\zeta_{D,\gamma}$ has increased, we repeat the loop. The process stops when the cardinality of $\zeta_{D,\gamma}$ stabilizes. This yields the set $\zeta_{D,\gamma}$.
Checking the sufficient condition for the support $D$ is straightforward. We iterate over the row indexed by $\gamma$ and verify the condition from Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$.
The necessary condition is checked as follows. We describe the $i$-th step of the verification. We examine the row indexed by $\delta_{i}$ (recall that $\delta_{1}=\gamma$). If we encounter a non-zero element $\lambda([e_{\delta_{i}},e_{\varepsilon}])$ such that $\varepsilon\notin\zeta_{D,\gamma}$ and $\varepsilon\neq\varepsilon_{i-1}$ (where $\varepsilon_{0}$ is undefined), we inspect the column indexed by $\varepsilon$, starting from row $\gamma$ upwards. If there are no other non-zero elements in this segment of the column, the necessary condition is not satisfied. If this is not the case, but exactly one such element $\lambda([e_{\delta_{i}},e_{\varepsilon}])$ with $\varepsilon\notin\zeta_{D,\gamma}$ is found in the row (ignoring a possible $\lambda([e_{\delta_{i}},e_{\varepsilon_{i-1}}])$), and if, in the corresponding segment of the column, there is exactly one other non-zero element of the form $\lambda([e_{\delta},e_{\varepsilon}])$, then we set $\varepsilon_{i}=\varepsilon$ and $\delta_{i+1}=\delta$. We then proceed to check the condition for the next column, indexed by $\delta_{i+1}$.
In practice, our program checks the necessary condition only up to the second step. If the condition has not been violated by that point, the program considers the check passed. This simplification does not significantly affect the final result.
@@
The results of the check are as follows: 878 supports satisfy the sufficient condition, while 911 satisfy the necessary condition. Therefore, 33 supports remain to be checked manually. They are listed below:
@@
Linear forms with support 5 do not satisfy the necessary condition, but checking this requires three steps, whereas our algorithm checks only the first two steps (see above). It is clear that linear forms with supports numbered 2, 6, 12, 14, 16, 18, 19, 20 and 21 do not satisfy condition (1); that is, such forms do not belong to $S$. This follows immediately from Table 1, similar to Example Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$. In the remaining cases, it is necessary to write out the coordinates of the infinitesimal action explicitly. It turns out that the linear forms with supports numbered 25 and 33 lie in $S$ without any restrictions on their coordinates. Linear forms with supports numbered 30, 31, and 32 lie in $S$ if and only if $\lambda_{\alpha_{1}+\alpha_{2}}\lambda^{2}_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}=\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}\lambda^{2}_{\alpha_{2}+\alpha_{3}+\alpha_{4}}$. Thus, we obtain 883 supports whose corresponding forms belong to $S$, three of which require an additional condition on the coordinates. This completes the proof of Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$.
Example 2.5. Consider a form $\lambda$ with support 33. We will show that $\lambda\in S$. Let $\mu\in\mathfrak{n}.\lambda,~\mu=x.\lambda,~x=\sum\nolimits_{\alpha\in\Phi^{+}}x_{\alpha}e_{\alpha},~\lambda=\sum\nolimits_{\alpha\in\Phi^{+}}\lambda_{\alpha}e_{\alpha}^{*},~\mu=\sum\nolimits_{\alpha\in\Phi^{+}}\mu_{\alpha}e_{\alpha}^{*}$. Let us write down explicitly the coordinates of interest. We denote the structure constants in the relation $Ce_{\alpha+\beta}=[e_{\alpha},e_{\beta}]$ by $C_{\alpha,\beta}$. Their values can be found in [KP23]. Then
@@
Let us also consider support number 25. For all roots except $\alpha_{2}$, condition (1) obviously holds. The coordinate $\mu_{\alpha_{2}}$ equals $x_{\alpha_{3}}\lambda_{\alpha_{2}+\alpha_{3}}$. Therefore, if $x_{\alpha_{3}}$ can be expressed as a linear combination of $\{\mu_{\beta}\}_{\beta\succ\alpha_{2}}$, then $\alpha_{2}$ satisfies condition (1). This is indeed the case:
@@
Finally, consider forms with supports 30, 31, and 32. Suppose the support contains the roots $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4}$. Let us check whether condition (1) holds for the root $\alpha_{2}$. Note that
@@
We see that if $\beta\succ\alpha_{2}$ and the variable $x_{\alpha_{3}+\alpha_{4}}$ appears in $\mu_{\beta}$, then $\beta=\alpha_{2}+\alpha_{3}$. Similarly, if $\beta\succ\alpha_{2}$ and $x_{\alpha_{1}}$ appears in $\mu_{\beta}$, then $\beta=\alpha_{2}+2\alpha_{3}$. Hence, the coordinate $\mu_{\alpha_{2}}$ must be a linear combination of $\mu_{\alpha_{2}+2\alpha_{3}}$ and $\mu_{\alpha_{2}+\alpha_{3}}$. It is easy to verify that this holds if and only if $\lambda_{\alpha_{1}+\alpha_{2}}\lambda^{2}_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}=\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}\lambda^{2}_{\alpha_{2}+\alpha_{3}+\alpha_{4}}$.
Now, let us check condition (1) for the root $\alpha_{4}$. Note that
@@
By an argument analogous to the one above, we find that the coordinate $\mu_{\alpha_{4}}$ must be a linear combination of $\mu_{\alpha_{1}}$ and $\mu_{\alpha_{3}+\alpha_{4}}$. It turns out that this condition is also equivalent to $\lambda_{\alpha_{1}+\alpha_{2}}\lambda^{2}_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}=\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}\lambda^{2}_{\alpha_{2}+\alpha_{3}+\alpha_{4}}$.